OpenStudy (anonymous):

Solve for t: 3t^2 - t =2 A. t= 3,2 B. t= 1/2, 1 C. t= -2/3,1 D. t= -1/2,1

OpenStudy (zzr0ck3r):

we want the form \[at^2+bt+c=0\] we have \[3t^2-t=2\\\downarrow\\3t^2-t-2=0\] we want two numbers that multiple to a*c=3*-2=-6 but we want these same two numbers to add to b = -1 what are two such numbers?

OpenStudy (anonymous):

I'm not understanding how you find these numbers.

OpenStudy (anonymous):

transpose 2 to the left side, then factor the equation.

OpenStudy (anonymous):

3t^2-t-2=0

OpenStudy (anonymous):

the factor it using quadratic equation or any method that's easier for you. quadratic equation: